OpenStudy (anonymous):

(iv) There is no perfect cube which ends with 8.true or false

OpenStudy (anonymous):

false

OpenStudy (anonymous):

this may be stupid but i thought 8 was a perfect cube since 2^3 =8

OpenStudy (anonymous):

False 2^3 = 8 12^3 = 1728